Exercise 11-1 Payback period computation; uneven cash flows LO P1

Beyer Company is considering the purchase of an asset for $290,000. It is expected to produce the following net cash flows. The cash flows occur evenly within each year. Year 1 Year 2 Year 3 Year 4 Year 5 Total Net cash flows $ 70,000 $ 40,000 $ 70,000 $ 200,000 $ 20,000 $ 400,000 Compute the payback period for this investment.

(Cumulative net cash outflows must be entered with a minus sign. Round your Payback Period answer to 2 decimal place.)
Answer #1

Payback period refers to the time in which initial investment in the project is recovered.

Payback period = A + B/C

Where,

A = years of full recovery

B = Uncovered cost at the beginning of the last year

C = Cash flows in the following year

Payback period = Cost of machine/Annual cash inflow

year Cash inflow cumulative cash inflow
1 70,000 70,000
2 40,000 110,000
3 70,000 180,000
4 200,000 380,000
5 20,000 400,000

Payback period = 3 + 110,000/200,000

= 3 + 0.55

= 3.55 years

Please do comment if you have any query, thanks.